# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@
Network interference, message-passing model, experimental design, network data, high-dimensional data
@@ 1 Introduction
Randomized Control Trials (RCTs) are the gold standard for evaluating intervention effectiveness, such as assessing public campaigns to promote voting within communities (Imbens and Rubin, 2015). However, classic RCT methods often overlook the complex dynamics of belief adoption within social networks. In these networks, experimental units—individual voters, for instance—do not make decisions in isolation but are influenced by peers, family, and societal norms. A study of the 2010 US congressional elections by Bond et al., (2012) demonstrated that encouraging individuals to vote influenced not only their own behavior but also that of their friends and those beyond their immediate social circles.
This phenomenon, known as network interference, violates a fundamental RCT assumption—the Stable Unit Treatment Value Assumption (SUTVA)—which requires that the treatment status of one unit does not affect the outcomes of other units. This violation precludes clear separation of treatment and control groups, a challenge prevalent in modern social science contexts (Imbens, 2024). This raises a central question across scientific fields: “How can we efficiently estimate and validate the causal effect of an intervention within a network of interacting units?” (Eckles et al., 2016, Cai et al., 2015, Abaluck et al., 2022, Ogburn et al., 2024).
To articulate the challenges in the estimation and validation of causal effects, we first introduce a formal notation and visualization. Consider a randomized experiment conducted over time stamps $t=0,1,\ldots,T$, involving a population of $N$ units, indexed by $i=1,\ldots,N$. At each time $t$, unit $i$ receives a treatment (e.g., exposure to the voting campaign) according to a random variable $W^{i}_{t}$. In the voting example, these treatment variables are binary: 1 indicates assignment to the treatment group, and 0 indicates assignment to the control group. [^1] Our theoretical results apply to the more general case when the treatments can also be continuous-valued.
@@
With $\bm{w}{}$ being a specific realization of $\bm{W}{}$, let $\bm{Y}{}(\bm{W}{}=\bm{w}{})$ denote an $N\times(T+1)$ matrix that collectively represents the observed outcomes under the treatment allocation $\bm{w}{}$. For example, if per-unit treatment probability at all time periods is equal to $p$, $\text{CFE}_{t}{(\bm{W}{})}$ evolves as a stochastic function of time and $p$, where the randomness arises from the treatment allocation $\bm{W}{}$. Figure 1 visualizes this evolution as a two-dimensional surface parameterized by time and $p$, where each contour (at fixed $p$) represents one possible dynamic trajectory. [^3] Note that treatment probabilities can vary over time, making the space of possible counterfactual evolutions $\text{CFE}_{t}{(\bm{W}{}=\bm{w}{})}$ considerably more complex than the simplified visualization shown in Figure 1.
@@
Figure 1: Evolution of outcomes sample mean (z-axis) with respect to time $t$ and treatment probability $p$. Red and blue contours highlight the counterfactual evolutions at treatment probabilities $p=0.25$ and $p=0.75$, respectively. The magenta line represents the equilibrium state, where the treatment effect has stabilized.
@@
Our objective is to estimate counterfactual evolutions $\text{CFE}_{t}{(\bm{w}{}^{\prime})}$ under alternative treatment allocations $\bm{w}{}^{\prime}\neq\bm{w}{}$ and validate the accuracy of these estimations. This enables estimation of treatment effects over time through comparing counterfactual evolutions under different treatment conditions—for example, in the voting study, contrasting evolution under an intensive campaign against the status quo control. However, this objective faces a fundamental challenge: we can observe the system only under a single scenario, reflecting “the fundamental problem of causal inference” (Holland, 1986). Specifically, with treatment probability fixed at $p$, we observe only the corresponding contour in the surface shown in Figure 1. This limitation constrains both our ability to estimate counterfactual evolutions and validate such estimates, as we cannot directly observe outcomes under alternative scenarios.
@@
To address the estimation challenge, we develop our approach by drawing an analogy between two systems: a network of units exchanging causal effects and a molecular system exchanging energy. In molecular systems, energy perturbations trigger redistribution, according to invariant physical laws governing molecular interactions, until a new equilibrium is reached. Similarly, in social networks, targeted interventions propagate through network connections following consistent social and behavioral principles until the system stabilizes. This analogy highlights a crucial insight: the temporal invariance of underlying mechanisms. Just as the equations governing molecular energy redistribution maintain their structural form, the mechanisms driving the propagation of intervention effects preserve their fundamental structure over time. This structural invariance presents a key opportunity: by observing the system’s evolution across different time periods, we can employ machine learning tools to learn the mathematical functions that characterize how intervention effects propagate through the network and aggregate into counterfactual evolutions [^4] This structural invariance differs from stationarity, as we allow the underlying mechanisms to exhibit temporal variations while maintaining their basic mathematical form..
@@
To address the validation challenge, a natural candidate is cross-validation—a cornerstone of statistical learning (Hastie et al., 2009) whose theoretical foundations (Stone, 1974, Allen, 1974) build upon Fisher’s fundamental insight about the necessity of randomization for independent validation (Fisher, 1935). However, cross-validation’s application in causal inference has been notably limited by data scarcity. Existing analyses typically focus on observations at equilibrium reached at large $t$, depicted by the “equilibrium” curve in Figure 1 (Basse et al., 2019, Jackson et al., 2020, Li and Wager, 2022b ), providing insufficient samples for reliable validation. Even recent approaches that leverage temporal pre-equilibrium data (Shirani and Bayati, 2024, Bayati et al., 2024) provide only order $T$ summary statistics. This represents a significant constraint, given that extending experiment duration is often prohibitively costly (Holtz et al., 2020, Cooprider and Nassiri, 2023, Xiong et al., 2024a ). To overcome this limitation, we introduce a new distribution-preserving network bootstrap method (DPNB) that generates sufficient data to implement cross-validation, enabling rigorous validation of our estimation results.
@@
By combining the temporal invariance principle with DPNB, we create a comprehensive framework that addresses both estimation and validation challenges in settings with unobserved network interference. Specifically, this work makes four main contributions. First, we extend the Causal Message-Passing framework (Shirani and Bayati, 2024) by incorporating unit- and network-level heterogeneities, deriving more general equations for outcome evolution. Second, we develop non-asymptotic versions of these dynamic equations using techniques from Li and Wei, (2022), leading to our DPNB method for generating and analyzing multiple subpopulations from the experimental population. This approach yields sufficient observations at each time period for accurate estimation through machine learning techniques.
Third, we combine temporal data with our DPNB method to develop a cross-validation framework for counterfactual estimation, implementing data-driven model selection procedures and preprocessing for complex time-trends to improve accuracy. Our DPNB method may have applications beyond causal inference, particularly in network sampling problems, where it offers an approach for generating representative subsamples from networked populations without requiring prior knowledge of network structure.
@@
Our fourth contribution addresses a fundamental challenge in causal inference: the rigorous examination of estimation methods when ground-truth counterfactuals are typically unobservable. We develop a comprehensive benchmark toolbox with six diverse experimental environments that capture different forms of network interference and temporal dynamics. These environments encompass applications ranging from social network influence to data center load balancing and ride-sharing systems. A notable highlight is our social network environment, which simulates a social media platform through thousands of AI agents interacting with their local network connections via content feeds. Each environment is carefully constructed to maintain known ground truth values while incorporating realistic complexities such as time-varying trends and heterogeneous network effects. We use this toolbox to extensively evaluate our method against existing benchmarks. Importantly, we are making these environments publicly available to facilitate standardized testing of future methodologies. [^5] Code and environments available at https://github.com/CausalMP/CausalMP.git. This contribution advances the field by providing researchers with a much-needed framework for systematic evaluations of causal inference methods under general network interference (Basse and Airoldi, 2018, Sävje et al., 2021, Arkhangelsky and Imbens, 2023, Imbens, 2024).
@@
The rest of the paper is organized as follows. Section 2 reviews related literature on causal inference under network interference, approximate message passing, and network sampling. Section 3 presents our generalization of the Causal Message-Passing framework (Shirani and Bayati, 2024). Section 4 introduces our distribution-preserving network bootstrap method for generating sufficient data to learn invariant aggregate dynamics. Section 5 presents our cross-validation framework for counterfactual estimation. Sections 6 and 7 describe our benchmark toolbox and empirical validation results. Section 8 concludes with discussion and future directions. Technical proofs and supplementary results appear in the appendices.
@@ 2 Related Literature
Recent literature on causal inference under network interference has evolved along several key dimensions. One common approach addresses settings with known network clusters (Sobel, 2006, Rosenbaum, 2007, Hudgens and Halloran, 2012, Tchetgen and VanderWeele, 2012, Auerbach and Tabord-Meehan, 2021), where researchers randomly assign entire clusters to different treatment intensities. This cluster-based design helps contain interference within cluster boundaries, though it requires prior knowledge of network clusters. Another significant development employs “exposure mappings” (Manski, 2013, Aronow and Samii, 2017, Leung, 2022, Harshaw et al., 2023, Sävje, 2024), which quantify how a unit’s treatment effect varies based on its neighbors’ treatment status. For scenarios with unknown interference patterns, researchers have proposed alternative strategies using historical data or staggered treatment rollouts (Yu et al., 2022, Cortez et al., 2022). A particularly promising direction has emerged through application-specific approaches that develop customized interference models for distinct contexts such as marketplace dynamics (Bajari et al., 2021, Holtz et al., 2020, Wager and Xu, 2021, Munro et al., 2021, Johari et al., 2022, Bright et al., 2022). These domain-specific solutions offer tailored frameworks for managing interference in their respective experimental settings. Our work contributes to this literature by developing methods that do not require knowledge of the interference structure or domain-specific context while still capturing complex network effects across diverse applications.
@@
Our methodology is supported by rigorous theoretical results derived from Approximate Message-Passing (AMP), whose foundations trace back to Thouless et al., (1977), Kabashima, (2003) and Donoho et al., (2009) and were formally established by Bolthausen, (2014), Bayati and Montanari, (2011) and a large body of recent literature (Javanmard and Montanari, 2013, Bayati et al., 2015, Berthier et al., 2020, Chen and Lam, 2020, Zhong et al., 2021, Wang et al., 2022, Dudeja et al., 2023, Rush and Venkataramanan, 2018, Li and Wei, 2022). At its core, we employ the TAP equation and cavity method—a mathematical technique that characterizes system behavior by analyzing responses to carefully introduced perturbations (Mezard et al., 1986, Mezard and Montanari, 2009). AMP traditionally addresses high-dimensional estimation problems through iterative systems with known mixing matrices. Our approach is different and builds on the work of Shirani and Bayati, (2024); specifically, they assume the mixing matrix is unknown and develop techniques similar to those developed in the AMP literature to analyze observed outcomes and infer causal effects.
@@
Our work on distribution-preserving network bootstrap relates to a rich body of research on network sampling and statistical inference in networked settings. The challenge of generating representative samples from network data has been central to network analysis, due to the inherent dependencies in network structures (Wormald, 1999, Newman, 2018). Network inference methods have evolved from basic configuration models (Bollobás, 1980) to sophisticated approaches handling non-independent observations (Snijders and Borgatti, 1999, Bickel and Chen, 2009) and robust bootstrapping techniques (Bhattacharyya and Bickel, 2015, Green and Shalizi, 2022). Recent advances encompass null models for complex network structures (Karrer and Newman, 2011, Bianconi, 2018) and machine learning approaches using graph neural networks (Kipf and Welling, 2017, Hamilton, 2020). Our approach differs fundamentally from this literature as it operates solely on node-level outcomes observed over time, without requiring access to the underlying network structure.
3 General Counterfactual Estimation
@@
Estimating counterfactual evolution $\left\{\text{CFE}_{t}{(\bm{w}{}^{\prime})}\right\}_{t=0}^{T}$ based on the observed outcomes $\bm{Y}{}(\bm{W}{}=\bm{w}{})$ enables addressing a broad range of causal questions, such as, “What if we had delivered the treatments according to $\bm{w}{}^{\prime}$ instead of $\bm{w}{}$?” For example, if outcomes are observed after treating 20% of potential voters, campaign organizers may want to explore how the population would have responded if 40% had received the campaign materials. In addition, by estimating the dynamics of the counterfactuals over time, we gain insights into how the treatment effect may strengthen or weaken as time progresses.
@@
The total treatment effect (TTE) provides a formal way to compare any two counterfactual scenarios. For two treatment allocations $\bm{w}{}^{\prime}$ and $\bm{w}{}^{\prime\prime}$, the TTE measures the difference in population average outcomes:
@@
The literature typically focuses on a special case: when all entries of $\bm{w}{}^{\prime\prime}$ equal one and all entries of $\bm{w}{}^{\prime}$ equal zero, with all outcomes at equilibrium (Yu et al., 2022, Candogan et al., 2023, Ni et al., 2023, Ugander and Yin, 2023). However, this extreme scenario of treating the entire population versus treating no one is often impractical. As we demonstrate in subsequent sections, our framework enables the estimation of TTEs between any two treatment allocations, providing decision-makers with more realistic comparisons (Muralidharan and Niehaus, 2017, Egger et al., 2022).

### v2

@@
Experimental design, network interference, cross validation, distribution-preserving bootstrap
@@ 1 Introduction
Randomized Controlled Trials (RCTs) are the gold standard for evaluating intervention effectiveness (Imbens and Rubin, 2015), such as assessing the impact of a public health initiative like deworming program to reduce parasitic infections in school-aged children. However, classic RCT methods often overlook the complex dynamics of disease transmission within communities. In these settings, experimental units, for example, individual children, are influenced by the infection status of their peers, family members, and broader community networks. A study of deworming interventions by Miguel and Kremer, (2004) demonstrated that treating some children for parasitic worms also influenced other children in neighboring schools through reduced environmental contamination and transmission pathways.
This phenomenon, known as network interference, violates the Stable Unit Treatment Value Assumption (SUTVA), a key principle of causal inference requiring that one unit’s treatment assignment not affect other units’ outcomes. When SUTVA is violated, the conceptual boundary between treatment and control groups dissolves, a methodological challenge found across many scientific domains. In public health, one person’s vaccination alters infection risk for others (Hudgens and Halloran, 2008); in economics, housing subsidies reshape neighborhood dynamics (Sobel, 2006); in healthcare, behavioral changes spread through social ties (Valente, 2012); in two-sided marketplaces, interventions targeting one side of the market create spillover effects on the other side (Johari et al., 2022); and in dynamic settings, current treatment effects depend on the history of past interventions. This pervasiveness raises a fundamental question: “How can we efficiently estimate and validate the causal effect of an intervention within a network of interacting units?” (Ugander et al., 2013, Eckles et al., 2016, Abaluck et al., 2022, Farias et al., 2022, 2023, Ogburn et al., 2024, Imbens, 2024, Johari et al., 2024).
Figure 1: Validating Counterfactual Estimations: (1) Network interventions evolve through a propagation phase, governed by invariant mathematical rules known as state evolution equations. (2) Distribution-preserving network bootstrap generates different samples that retain the statistical properties of the original experimental population. (3) Counterfactual cross-validation partitions the time horizon to train and validate estimation models. (4) Six experimental environments spanning diverse domains and network structures to test estimation methods.
To articulate the challenges in the estimation and validation of causal effects, we first introduce a formal notation and visualization. Consider a randomized experiment conducted over time stamps $t=0,1,\ldots,T$, involving a population of $N$ units (e.g., individual children in the deworming study), indexed by $i=1,\ldots,N$. At each time $t$, unit $i$ receives a treatment (e.g., the medication) according to a random variable $W^{i}_{t}$. In the deworming example, these treatment variables are binary: 1 indicates assignment to the treatment group, and 0 indicates assignment to the control group.
@@
With $\bm{w}{}$ being a specific realization of $\bm{W}{}$, let $\bm{Y}{}(\bm{W}{}=\bm{w}{})$ denote an $N\times(T+1)$ matrix that collectively represents the observed outcomes under the treatment allocation $\bm{w}{}$. For example, if per-unit treatment probability at all time periods is equal to $p$, the sample mean $\text{CFE}_{t}{(\bm{W}{})}$ is a random variable whose randomness arises from the treatment allocation $\bm{W}{}$ and potentially from inherent stochasticity in the outcomes themselves. However, for a large $N$, this sample mean converges to its expectation. This deterministic evolution is visualized as the smooth surface in Figure 2, where each contour (at a fixed $p$) represents the expected dynamic trajectory. [^3] Note that treatment probabilities can vary over time, making the space of possible counterfactual evolutions $\text{CFE}_{t}{(\bm{W}{}=\bm{w}{})}$ considerably more complex than the simplified visualization shown in Figure 2.
@@
Figure 2: Evolution of outcomes sample mean (z-axis) with respect to time $t$ and treatment probability pp. Red and blue contours highlight the counterfactual evolutions at treatment probabilities $p=0.25$ and $p=0.75$, respectively. The magenta line represents the equilibrium state, where the treatment effect has stabilized.
@@
Our objective is to estimate expected counterfactual evolutions $\text{CFE}_{t}{(\bm{w}{}^{\prime})}$ under alternative treatment allocations $\bm{w}{}^{\prime}\neq\bm{w}{}$ and validate the accuracy of these estimations. This enables estimation of treatment effects over time through comparing counterfactual evolutions under different treatment conditions; for example, we can contrast evolution under delivering the medication to everyone against the status quo control. However, this objective faces a fundamental challenge: we can observe the system only under a single scenario, reflecting “the fundamental problem of causal inference” (Holland, 1986). Specifically, with treatment probability fixed at $p$, we observe only the corresponding contour in the surface shown in Figure 2. This limitation constrains both our ability to estimate counterfactual evolutions and validate such estimates, as we cannot directly observe outcomes under alternative scenarios.
@@
To address the estimation challenge, we develop our approach by drawing an analogy between a network of units exchanging causal effects and a molecular system exchanging energy. In molecular systems, energy perturbations trigger redistribution according to invariant physical laws governing molecular interactions. Similarly, interventions propagate through network connections following consistent mechanical or behavioral principles. This analogy shows an important insight: the invariance of treatment propagation mechanisms. Just as the equations governing molecular energy redistribution maintain their structural form, the mechanisms driving the propagation of the intervention effects preserve their fundamental structure both over time and across treatment scenarios (Part 1 in Figure 1). This structural invariance gives us an opportunity: by observing system evolution over time, we can employ machine learning to learn the functions characterizing how intervention effects propagate and aggregate into counterfactual evolutions.
@@
To address the validation challenge, a natural candidate is cross-validation (Hastie et al., 2009) whose theoretical foundations (Stone, 1974, Allen, 1974) build upon Fisher’s insight about the necessity of randomization for validation (Fisher, 1935). However, the application of cross-validation in causal inference has been notably limited by data scarcity. Using observations at an equilibrium reached at large $t$, depicted by the “equilibrium” curve in Figure 2 (Basse et al., 2019, Jackson et al., 2020, Li and Wager, 2022b ), may not provide sufficient samples for reliable validation. Even using temporal pre-equilibrium data is insufficient, given that extending experiment duration is often prohibitively costly (Holtz et al., 2020, Cooprider and Nassiri, 2023, Xiong et al., 2024a ). To overcome this limitation, we introduce a new distribution-preserving network bootstrap method (DPNB) that generates sufficient data to enable systematic validation of our estimation methods (Parts 2 and 3 in Figure 1).
@@
By combining the invariant propagation principle with DPNB, this work advances our ability to address both estimation and validation challenges. Specifically, this work makes four main contributions. First, we extend the Causal Message-Passing (Shirani and Bayati, 2024) by incorporating unit- and network-level heterogeneities, capturing more general outcome evolution equations. Second, we develop non-asymptotic analyses of these dynamics using techniques from Li and Wei, (2022). This provides a principled way to study how two aspects of the interference network, its intensity and its heterogeneity, affect standard estimators such as difference-in-means. Specifically, we show that intensity drives the bias, while heterogeneity drives the variance. This analysis further leads to our DPNB method for generating and analyzing multiple subpopulations from the experimental population. This approach provides sufficient observations at each time period to enable accurate estimation through machine learning techniques. Third, we combine temporal data with our DPNB method to develop a cross-validation framework for counterfactual estimation that implements data-driven model selection. Applications of DPNB can go beyond causal inference, particularly in generating representative subsamples in various network problems without knowledge of the underlying network structure.
@@
Fourth, this work contributes to addressing a fundamental challenge in causal inference: rigorously evaluating estimation methods when ground-truth counterfactuals are typically unobservable. We develop a benchmark toolbox with six different experimental environments that capture different forms of network interference and temporal dynamics. These environments span applications from social network influence to data center load balancing and ride-sharing systems. A notable highlight is our social network environment, which simulates a social media platform using thousands of AI agents interacting through their local network connections via content feeds. Each environment is carefully constructed to maintain known ground truth values while incorporating realistic complexities such as time-varying trends and heterogeneous network effects. We use this toolbox to extensively evaluate our method and make these environments publicly available to facilitate standardized testing of future methodologies.
@@
The rest of the paper is organized as follows. Section 2 reviews related literature on causal inference under network interference, approximate message passing, and network sampling. Section 3 presents our generalization of the Causal Message-Passing framework (Shirani and Bayati, 2024). Section 4 develops our finite sample analysis of network interference dynamics, which leads to our distribution-preserving network bootstrap method for generating sufficient data to learn invariant aggregate dynamics. Section 5 presents our cross-validation framework for counterfactual estimation. Sections 6 and 7 describe our benchmark toolbox and empirical validation results. Section 8 concludes with discussion and future directions. Technical proofs and supplementary results appear in the appendices.
@@ 2 Related Literature
Recent literature on causal inference under network interference has evolved along several dimensions. One common approach addresses settings with known network clusters (Sobel, 2006, Rosenbaum, 2007, Hudgens and Halloran, 2008, Tchetgen and VanderWeele, 2012, Auerbach and Tabord-Meehan, 2021, Su and Ding, 2021, Viviano et al., 2023, Jia et al., 2024, Eichhorn et al., 2024), where researchers randomly assign entire clusters to different treatment intensities. This cluster-based design helps contain interference within cluster boundaries, though it requires prior knowledge of network clusters. Another significant development employs “exposure mappings” (Manski, 2013, Aronow and Samii, 2017, Leung, 2022, Harshaw et al., 2023, Sävje, 2024), which quantify how a unit’s treatment effect varies based on its neighbors’ treatment status.
For scenarios with unknown interference patterns, researchers have proposed alternative strategies using multi-period observations or design-based experiments (Yu et al., 2022, Cortez et al., 2022, Hu and Wager, 2022). A particularly promising direction has emerged through application-specific approaches that develop customized interference models for various contexts such as marketplace dynamics (Holtz et al., 2020, Bajari et al., 2021, Wager and Xu, 2021, Munro et al., 2021, Johari et al., 2022, Bright et al., 2022). These domain-specific solutions offer tailored frameworks for managing interference in their respective experimental settings (e.g., by leveraging the two-sided structure of the marketplace or assuming interference is mediated by market price). Our work contributes to this literature by developing methods that do not require knowledge of the interference structure or domain-specific context while still capturing complex network effects across different applications.
@@
Our methodology is supported by rigorous theoretical results derived from Approximate Message-Passing (AMP), whose foundations trace back to Thouless et al., (1977), Mézard et al., (1987), Kabashima, (2003), Mezard and Montanari, (2009), Donoho et al., (2009) and were formally established by Bolthausen, (2014), Bayati and Montanari, (2011) and a large body of recent literature (Javanmard and Montanari, 2013, Bayati et al., 2015, Berthier et al., 2020, Chen and Lam, 2020, Zhong et al., 2021, Wang et al., 2022, Dudeja et al., 2023, Rush and Venkataramanan, 2018, Li and Wei, 2022). AMP traditionally addresses high-dimensional estimation problems through iterative updates with known matrices. Our approach is different and builds on the work of Shirani and Bayati, (2024); specifically, we assume the matrix is unknown and develop techniques to analyze observed outcomes and infer causal effects.
@@
Our work on distribution-preserving network bootstrap relates to a rich body of research on network sampling and statistical inference in networked settings. The challenge of generating representative samples from network data has been important in network analysis, due to the inherent dependencies in network structures (Wormald, 1999, Newman, 2018). Network inference methods have evolved from basic configuration models (Bollobás, 1980) to sophisticated approaches handling non-independent observations (Snijders and Borgatti, 1999, Bickel and Chen, 2009) and robust bootstrapping techniques (Bhattacharyya and Bickel, 2015, Green and Shalizi, 2022). Another line of research focuses on the efficient algorithmic generation of random graphs that satisfy specific structural constraints, which is helpful for bootstrapping or creating null models (Milo et al., 2002). A central problem in this area is generating graphs with a fixed degree sequence. Classic approaches, such as Markov chain Monte Carlo methods (Jerrum and Sinclair, 1989) or switch-based algorithms (McKay and Wormald, 1990), can produce asymptotically uniform samples but are often computationally intensive, with runtimes that are high-degree polynomials. To address this, faster sequential construction algorithms have been developed, including those based on sequential importance sampling (Chen et al., 2005, Blitzstein and Diaconis, 2010) and other dynamic generation techniques (Steger and Wormald, 1999, Kim and Vu, 2004, Bayati et al., 2010, 2018). More recent advances also encompass general null models for complex network structures (Karrer and Newman, 2011, Bianconi, 2018) and machine learning approaches using graph neural networks (Kipf and Welling, 2017, Hamilton, 2020). Our approach differs fundamentally from this literature as it operates solely on node-level outcomes observed over time, without requiring access to the underlying network structure.
3 General Counterfactual Estimation Problem
@@
Estimating counterfactual evolution $\left\{\text{CFE}_{t}{(\bm{w}{}^{\prime})}\right\}_{t=0}^{T}$ based on the observed outcomes $\bm{Y}{}(\bm{W}{}=\bm{w}{})$ enables addressing a broad range of causal questions, such as, “What if we had delivered the treatments according to $\bm{w}{}^{\prime}$ instead of $\bm{w}{}$?” For example, delivering the deworming medication to 20% of children, we may want to explore how the population would have responded if 40% had received the medication. In addition, by estimating the dynamics of the counterfactuals over time, we gain insights into how the treatment effect may strengthen or weaken as time progresses.
@@
The total treatment effect (TTE) [^4] Also known as the global treatment effect (GTE). provides a formal way to compare any two counterfactual scenarios. For two treatment allocations $\bm{w}{}^{\prime}$ and $\bm{w}{}^{\prime\prime}$, the TTE measures the difference in population average outcomes:
@@
The literature typically focuses on a special case: when all entries of $\bm{w}{}^{\prime\prime}$ equal one and all entries of $\bm{w}{}^{\prime}$ equal zero, with all outcomes at equilibrium (Candogan et al., 2023, Ni et al., 2023, Ugander and Yin, 2023). However, this scenario of treating the entire population versus treating no one may be impractical in certain settings. Our goal is to estimate any desired counterfactual evolution $\text{CFE}_{t}{(\bm{w}{}^{\prime})}$ over the experimental horizon $t=0,1,\ldots,T$, as specified by Eq. (1). Such estimates enable the calculation of TTEs between arbitrary treatment levels and provide decision-makers with more realistic comparisons (Muralidharan and Niehaus, 2017, Egger et al., 2022).